OpenStudy (anonymous):

simplify (3/4+4/5i)-(1/2-3/10i)

hartnn (hartnn):

combine real and imaginary parts. whats 3/4 -1/2 =...? 4/5 - (-3/10) =... ?

OpenStudy (anonymous):

1/4 and 11/10

hartnn (hartnn):

thats absolutely correct! so its \(\large \dfrac{1}{4} + \dfrac{11i}{10}\) good! :)

OpenStudy (anonymous):

Thank you ( :

hartnn (hartnn):

welcome ^_^